### v1

# Hip Fracture Patient Pathways and Agent-based Modelling

Abstract
Increased healthcare demand is significantly straining European services. Digital solutions including advanced modelling techniques offer a promising solution to optimising patient flow without impacting day-to-day healthcare provision. In this work we outline an ongoing project that aims to optimise healthcare resources using agent-based simulations.
Index Terms:
agent-based simulation; patient flow; hip fracture.I Introduction
Several key factors contribute to recurring overcrowding in Irish hospitals, many of which reflect deep-rooted systemic issues. A major issue driving increased healthcare demand, is the complex medical needs and increased ED attendance frequency required for older patients (patients aged years). Ireland’s rapidly ageing population directly correlates to increased patient demand [WALSH2021] and the European Union (EU) notes a similar trend [EUROSTAT2020, EC2018]. Data from 2022 paints a stark picture of the severity of the problem. In 2022 over 70% of Irish emergency departments exceeded their maximum capacity, leaving staff and resources perilously stretched [HIQA2022c]. Even more concerning, 80% of these hospitals reported significant staffing shortages, further compounding the difficulties in managing patient care effectively [CULLEN2022, POWER2022a]. The shortage of trained healthcare workers has been a persistent issue across the EU [LOOI2024]. We’re now faced with healthcare environments increasingly unable to meet demand [HSE2022].
In response to these challenges, Ireland’s regulatory body issued a series of recommendations emphasising the importance of effective workforce planning [HIQA2022b]. More recently they have demonstrated a strong appetite within Ireland for healthcare digital technologies [WHELAN24]. Digital technologies offer a promising solution but these efforts must be paired with forward-thinking strategies to prepare for future challenges. Without decisive action, the overcrowding crisis will persist, with dire societal consequences.
II Artificial Intelligence
Digital technology is revolutionising healthcare by offering new insights through data analysis. Automation reduces the need for manual coordination between departments, reduces the administrative burden of healthcare workers [ABBING2016], enabling them to focus on patient care, thus increasing the overall efficiency of the hospital [KELLY2022]. Machine learning (ML) is a sub-field of Artifical Intelligence (AI) that addresses the methods used by AI to perform specific tasks [EP2022]. ML algorithms have demonstrated abilities in predictive analytics and resource optimisation which can be used to improve operational efficiency [RATHORE2021b]. One of the most significant benefits of ML for healthcare applications is its ability to leverage existing data beyond the original data collection purposes. In Figure 1 we illustrate how existing data can be assessed to identify appropriate ML models providing actionable insights that can be used to optimise workflows.
The integration of AI/ML has long-term financial benefits for healthcare systems. One example is fewer instances of under-staffing, avoiding the financial strain associated with agency based staff, which cost Ireland in excess of €640 million in 2023 alone [WHITE2024].
While digital solutions offer the potential for transformative change in healthcare provision it is not without limitations. One limitation is the lack of transparency in algorithm outputs and enhancing explainability for algorithm outputs remains an active research area. The immense societal benefits of digital health has been recognised at a European level [EC2021]. For hospitals already facing strain due to rising patient numbers and limited resources, these technologies are fast becoming a critical requirement for ensuring that patient care is delivered efficiently without compromising quality. The benefits extend to long-term cost reductions, better resource management, and improvements in patient outcomes, all of which contribute to a more sustainable and effective healthcare system [BUIJS2024].
The question is no longer whether we will implement these technologies, but rather how we can do so effectively and responsibly. In this ongoing collaborative research project, we are investigating the integration of ML technologies into a Model 4 hospital in the mid-west of Ireland. This initiative aims to explore applications of ML in areas such as predictive diagnostics, patient flow optimisation, and resource management, potentially revolutionising healthcare delivery. Through this research, we seek not only to enhance patient care and hospital efficiency at a local level but also to establish a framework that could scale nationally, transforming the broader healthcare landscape. Ultimately, the project aspires to position Ireland at the forefront of healthcare innovation, with the potential for these technologies to be adopted on a global scale. Figure 2 details the data sources and ML tools of interest to this research, which is discussed in more detail later.
III Hip fractures: a case study
The ageing population of Europe has been identified as a significant contributor to overcrowding. Considering the correlation between age and fragility it is not surprising that older people (+65) are at a higher risk of experiencing complex medical fractures [WALSH2021]. Literature [KANIS2021] indicates the European cost of fracture services to be in excess of €56 billion (2019) with €290 million attributed to the Irish exchequer. Considering patient flow for a specific sub-set of patients, in this case hip fracture patients, we aim to identify specific causes of poor patient flow enabling targeted resource allocation that impacts both subset and larger population groups. Here we detail our ongoing research into how digital technologies can be used to improve system operations across an acute care setting for this patient group.
Mid-west Ireland has a population of approximately 500,000 with healthcare provided by the UL Hospital Group (ULHG). The ULHG is comprised of six facilities with University Hospital Limerick (UHL) as the regions Model 4 facility operating as the centralised location for critical care [HSE2023b]. UHL monitor patients from admission in the ED through the wider acute care setting (e.g., surgical and/or medical procedures) up to patient discharge. UHL also liaise with the National Office of Clinical Audit (NOCA). NOCA maintains the Irish Hip Fracture Database (IHFD) which collates hip-fracture patient information nationally. The purpose of the IHFD is to assess individual hospital compliance against seven key performance indicators (KPIs) including patient time to ward, patient time to the operating theatre [KELLY2022]. While the overall purpose of the IHFD is to assure quality of healthcare provision, the richness of the data within the IHFD can be exploited far beyond this task. Forecasts predict that hip fracture hospitalisations in Ireland could increase threefold by 2046 [KELLY2018], which would have a significant impact on HSE resourcing. In 2019 the +65 cohort increased by 1.2% while a 4.2% increase in hip fracture cases was reported [JONES2020].
In Figure 2 we illustrated the specific datasets and tools of interest to this work. The outer circle of Figure 2 shows datasets that can be used to provide additional context of value and relevance to hip-fracture patients. Immigration, due to global challenges, has significantly increased Ireland’s population in recent years with the population increasing by over 98,000 in the year ending April 2024 [CSO2024a] alone. The share of elderly ( years) immigrants has grown in the EU to account for approximately 21% of the immigrant population and approximately 6% are over 75 years [OECD2023]. This has obvious implications for age-related illnesses such as hip fractures where the mean age of fractures is 83 and 84 years for men and women, respectively. Analysing national census databases [CSO2024b] enables us to statistically interrogate how national and regional population demographics may affect service demand. Similarly, patient information captured by UHL can provide insights on several levels. The IHFD will provide contextual information related to the average patient age, and the frequency of hip-fractures nationally and regionally. Most patients suffering from hip-fractures arrive in acute care environments are posited to arrive via ambulance services but, to our knowledge, this relationship has not been confirmed. Using National Ambulance Service (NAS) records to assess the frequency of NAS hip-fracture call-outs enable us to better understand the resource draw on NAS for this patient group. Previous literature [STANLEY2023] indicated an association between national weather warnings and hip fractures. However, national weather warnings typically do not indicate regional specific affects and the type of weather (e.g., rain, ice, snow etc) was not addressed. International work [YEUNG2020] demonstrates a strong correlation between night-freezing weather events and fall-related injuries.
The middle circle of Figure 2 shows the machine learning algorithms identified as potential tools for analysing data. Time-series analysis can be employed to predict seasonal hip-fracture service needs, the average patient length of stay and the average time spent EDs and/or surgical units. We previously (Section I) outlined how explainability in AI/ML healthcare is a core requirement. In this research we will investigate multiple types of ML algorithms (see Figure 1) and assess their performance in terms of accuracy. One of the KPIs in our research will centre around explainable artificial intelligence comparing so-called ‘black-box’ algorithm performance with more interpretable ‘white- or glass- box’ methods. More advanced simulation methods such as agent-based models, which simulate interaction events, can be used to not only identify the most common patient pathways but also indicate pathway bottlenecks. Agent-based models enable us to interrogate multiple ‘what-if’ scenarios without impacting day-to-day functioning of acute environments. Test cases of interest to this work include scenarios such as: 1. What is the impact of ageing on current service demand and how will this change in the future based on population changes? 2. What is the optimal staffing resources required to maximise ED service? 3. How do optimal resource requirements change for specific patient groups?
The inner circle of Figure 2 illustrates our expected end goal. Our suite of tools will be used to create a dynamic dashboard illustrating in an easily interpretable manner key information related to this patient subgroup. As this work is ongoing the dashboard display has not yet been fully conceptualised but is likely to include information such as: current number of patients, model prediction of future patients, and the average patient wait time in ED.
IV Impact
In the context of the IHFD it is interesting to note that data collection frequently occurs post-intervention which limits the ability of healthcare workers to effect change in a timely manner. Using digital platforms like that illustrated in Figure 2 to automate both data collection and data analysis, opining the possibility of real-time information that can be used to provide an improved patient experience. Further this type of system can be integrated with existing hospital data collection systems reducing the need for manual, repetitive administration. Accessing data in this manner reduces treatment delays benefiting patients [KELLY2022]. Accurate ML models could be used to support operational decision-making around workforce planning and resource allocation. Armed with these predictions, hospitals can preemptively adjust staffing levels, prepare operating theatres, and ensure that enough beds are available to meet patient demand. This type of foresight allows hospitals to avoid being overwhelmed, optimises resource use and prevents operational inefficiencies. By better managing both human and material resources, hospitals can operate more cost-effectively, benefiting not only the institution but also the broader healthcare system in Ireland. The integration of datasets from diverse institutions enhances communication between multiple parties. The connectivity of a dynamic dashboard created from disparate sources can act as a starting point for larger interoperable systems extending beyond acute care environments to incorporate national and regional population data, as outlined here. Future work includes aims to integrate acute- and community- care (e.g. general practitioners) systems providing a more holistic healthcare system for Irish citizens.
V Conclusion
In summary, the integration of digital technology, AI, and ML in healthcare represents a transformative opportunity to address the pressing challenges faced by emergency departments, particularly in the context of the IHFD. The continuous surges in patient volume, especially in emergency settings, have led to significant overcrowding and strain on the healthcare system. By leveraging data-driven insights, hospitals can optimise resource allocation, improve patient flow, and enhance compliance with KPIs. The methods discussed here offer a unique insight into how acute care data collection in Ireland can be leveraged beyond it’s original purpose providing actionable insights that can radically improve workflows without negatively impacting daily acute care operations. These technologies not only facilitate the forecasting of patient admissions but also enhance the prioritisation of care for high-risk patients, ultimately leading to improved patient outcomes. Moreover, the application of AI and ML allows for the continuous optimisation of patient care pathways, ensuring that healthcare resources are utilised effectively and efficiently.
As healthcare systems strive to improve efficiency and patient care, the adoption of these digital technologies and analytical methods will be crucial. The potential for long-term cost savings, improved patient outcomes, and more effective resource management underscores the importance of embracing this technological evolution in healthcare. Ultimately, by harnessing the power of ML hospitals can not only meet current demands but also proactively prepare for future challenges, creating a more resilient and responsive healthcare system for all.
VI Acknowledgements
We gratefully acknowledge the expertise, guidance and support of our UHL colleagues Damien Ryan, Carrie Garavan and Fergal Cummins. This project is funded under the National Challenge Fund which was established under the government’s National Recovery and Resilience Plan (NRRP), funded by the EU’s Recovery and Resilience Facility. The fund is coordinated and administered by Taighde Éireann – Research Ireland grant 22/NCF/DR/11289.